find the solution of the system of linear equation
2x+4y=4
2x-4y=8

Found 2 solutions by oscargut, jim_thompson5910:
Answer by oscargut(2103)

2x+4y=4
2x-4y=8
Adding them:
4x = 12
x = 3
y = -1/2
Answer: x = 3 and y = 1/2
You can ask me more at: mthman@gmail.com
Thanks

Answer by jim_thompson5910(35256)

Add the equations to get

2x+4y=4
2x-4y=8
--------
4x+0y = 12



4x = 12

x = 12/4

x = 3


Now use this to find y

2x+4y=4

2(3) + 4y = 4

6 + 4y = 4

4y = 4 - 6

4y = -2

y = -2/4

y = -1/2

The solutions are x = 3 and y = -1/2